Determine: 

a) total number of atoms in 690.7 g of Pb

b) the number of moles of Si in a 312 g of sample

c) the mass of three trillion-atom (3.0x1012) sample of metallic gold

d) number of 41K atoms that are present in a piece of potassium block weighing 715g? 

(The percent isotope abundance of 41K is 6.73%?)


1
Expert's answer
2021-10-28T02:56:47-0400

a) 1mole of Pb= 207.2g = 6.022×10^23 atoms

Number of atoms =2×10^24

b)1mole of Si= 28

312g = 11.14moles

c) Imole of gold =197g= 6.022× 10^23

Mass= 9.8×10^-9

d) moles= 715/39= 18.33

Atoms= 1.1×10^25